# Reinforcement Learning for Reasoning in Small LLMs: What Works and What Doesn’t

### v1

@@ 4 Conclusion
Our study investigated enhancing the reasoning abilities of small LLMs using RL, focusing on the 1.5-billion-parameter DeepSeek-R1-Distill-Qwen-1.5B under strict constraints. Adapting the GRPO algorithm and a compact mathematical reasoning dataset, we conducted three experiments to assess behavior and performance under resource limitations. Our findings show small LLMs can achieve significant reasoning gains with minimal resources—e.g., AMC23 accuracy rising from 63% to 80% and AIME24 reaching 46.7%, surpassing o1-preview—at a cost of \$42 versus thousands for baselines. Open-RS variants averaged 53.0%–56.3% on benchmarks, demonstrating RL’s viability for small LLMs. Releasing our code and datasets, we provide a framework for lightweight, reasoning-capable models, despite challenges like optimization stability, laying a foundation for future work.
@@ Appendix A Related Work
A.1 Reasoning in Large Language Models
A substantial body of research has investigated methods to enhance the reasoning capabilities and factual accuracy of large language models (LLMs). Early approaches predominantly relied on prompting techniques to elicit structured reasoning. For instance, scratchpad-style prompting encourages models to break down problems into intermediate steps (Nye et al., 2021), while verification mechanisms assess the correctness of generated outputs (Cobbe et al., 2021). Chain-of-thought (CoT) prompting has emerged as a particularly effective strategy, leveraging demonstrations of step-by-step reasoning to improve performance on complex tasks (Wei et al., 2022; Kojima et al., 2022; Reynolds & McDonell, 2021). More recently, techniques such as intermediate self-reflection have been proposed to enable models to iteratively refine their reasoning processes (Shinn et al., 2023; Madaan et al., 2023).
In parallel, supervised fine-tuning (SFT) has been employed to embed reasoning capabilities directly into LLMs. Studies such as (Lewkowycz et al., 2022a) and (Rajani et al., 2019) demonstrate that fine-tuning on high-quality datasets can enhance problem-solving abilities. Notably, integrating CoT reasoning into SFT has shown significant promise; works like (Zelikman et al., 2022; Muennighoff et al., 2025; Ye et al., 2025) illustrate that fine-tuning with small, carefully curated datasets of CoT examples can yield substantial performance gains. However, these efforts have predominantly focused on large-scale LLMs, typically ranging from 7 billion to over 100 billion parameters. This reliance on massive models limits accessibility and practicality for resource-constrained settings, motivating the exploration of alternative approaches for smaller LLMs.
A.2 Reasoning with Reinforcement Learning
Reinforcement learning (RL) has emerged as a powerful paradigm for improving reasoning in LLMs, particularly for tackling complex, multi-step problems. Unlike SFT, which often optimizes for imitation of training data, RL enables models to learn from feedback, enhancing generalization to both in-distribution and out-of-distribution tasks (Chu et al., 2025; Yeo et al., 2025). Recent advancements underscore the efficacy of RL in this domain. For example, OpenAI (2024b) and DeepSeek-AI (2025) demonstrate that RL-based training can significantly boost reasoning performance, while Team (2025a) explores scaling laws for RL-driven LLMs. These studies highlight RL’s ability to refine decision-making processes by optimizing for task-specific rewards, such as correctness or logical coherence.
@@ B.1 Limitations
Third, the multilingual nature of the base model, DeepSeek-R1-Distill-Qwen-1.5B, introduced unintended language drift after 150–200 steps, despite efforts to enforce English-only outputs via prompts in Experiment 3. This limitation reflects a trade-off in using a pre-trained, multilingual foundation, which, while efficient, complicates monolingual optimization. Finally, our evaluation focused exclusively on mathematical reasoning benchmarks, leaving the generalizability of our approach to other domains—such as scientific reasoning or coding—unexplored. These constraints highlight the need for cautious interpretation of our results within the specified scope.
B.2 Discussion
@@
Our findings reveal a nuanced trade-off between efficiency and reasoning depth in small LLMs. The rapid performance gains observed in the first 50–100 steps across all experiments (Insight 1) suggest that small, high-quality datasets can effectively bootstrap reasoning capabilities, aligning with prior work on data efficiency in RL (Chu et al., 2025). However, the subsequent degradation underscores a sensitivity to over-optimization under fixed length constraints, a challenge also noted in larger models like DeepSeek-R1 (DeepSeek-AI, 2025). Experiment 2’s success with mixed difficulty levels (Insight 2) indicates that curriculum-like strategies could mitigate this. Meanwhile, the cosine reward’s stabilizing effect in Experiment 3 (Insight 3) suggests a promising direction for controlling reasoning verbosity, though it sacrifices peak accuracy compared to Experiment 2.
@@
Comparatively, our Open-RS variants achieved performance rivaling or exceeding state-of-the-art 1.5B models (e.g., DeepScaleR-1.5B-Preview) and even some 7B models, at a fraction of the cost and data volume. This efficiency challenges the prevailing reliance on massive datasets and computational resources in reasoning enhancement (OpenAI, 2024b; Luo et al., 2025), offering a scalable alternative for resource-constrained environments. However, the persistent multilingual drift and length limitations point to inherent challenges in adapting multilingual base models and optimizing for complex tasks within tight bounds.
B.3 Future Directions
@@
These limitations suggest several research avenues. First, extending training duration or employing multi-stage length schedules could address truncation issues, allowing the model to handle harder problems without compromising stability. Second, incorporating a lightweight language reward or monolingual pre-filtering of the base model might mitigate language drift, enhancing output consistency. Third, expanding the benchmark suite to include non-mathematical domains would test the generalizability of our approach, aligning with broader AGI goals. Finally, exploring hybrid methods—such as combining GRPO with search algorithms like MCTS (Feng et al., 2024)—could further deepen reasoning capacity without significantly increasing resource demands.
@@
In conclusion, our work demonstrates that RL-based fine-tuning can unlock substantial reasoning potential in small LLMs, even under stringent constraints. By identifying key trade-offs and offering practical insights, we pave the way for developing efficient, reasoning-capable models that balance performance and accessibility—a critical step toward democratizing advanced AI technologies.
@@ Appendix C Datasets
Detail datasets are used in Section 3.2.
  • AIME24  [^4] https://huggingface.co/datasets/AI-MO/aimo-validation-aime: 30 problems from the 2024 American Invitational Mathematics Examination, emphasizing advanced high-school-level reasoning.
@@
  • MATH-500 (Lightman et al., 2023b; Hendrycks et al., 2021): A subset of 500 problems from the MATH benchmark, sourced from various mathematics competitions and spanning algebra, calculus, and geometry.
  • Minerva (Lewkowycz et al., 2022b): 272 undergraduate-level problems across physics, biology, chemistry, economics, and other sciences, requiring quantitative reasoning (mathematics subset used).
  • OlympiadBench (He et al., 2024): 675 Olympiad-level problems in mathematics and physics, designed to challenge advanced reasoning abilities.
@@ Appendix D Baseline Models
The description of baseline models in Section 3.3.
@@
    • Llama-3.1-70B-Instruct (AI, 2024a): A 70B-parameter model optimized for instruction-following.
    • o1-preview (AI, 2024b): A high-performing reasoning model from OpenAI.
@@
    • Qwen-2.5-Math-7B-Instruct (Yang et al., 2024): An RL-trained model with a reward model for mathematical reasoning.
    • rStar-Math-7B (Guan et al., 2025): Uses Monte Carlo Tree Search (MCTS) for deep reasoning and self-evolution.
    • Eurus-2-7B-PRIME (Cui et al., 2025): Employs the PRIME method with online RL and process rewards.
    • Qwen2.5-7B-SimpleRL (Zeng et al., 2025): Applies Proximal Policy Optimization (PPO) with rewards based on final answers.
@@
    • DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025): The original untrained baseline.
    • DeepScaleR-1.5B-Preview (Luo et al., 2025): Fine-tuned with GRPO on 40,000 math problem-answer pairs across multiple RL stages.
    • Still-3-1.5B-Preview (Team, 2025b): RL-trained with a focus on slow thinking (e.g., tree search) using 30,000 curated math examples.
@@ Appendix E Hyperparameter Setup
Table 5 show parameters that used in training phase.

### v2

@@ 4 Conclusion
Our study investigated enhancing the reasoning abilities of small LLMs using RL, focusing on the 1.5-billion-parameter DeepSeek-R1-Distill-Qwen-1.5B under strict constraints. Adapting the GRPO algorithm and a compact mathematical reasoning dataset, we conducted three experiments to assess behavior and performance under resource limitations. Our findings show small LLMs can achieve significant reasoning gains with minimal resources - e.g., AMC23 accuracy rising from 63% to 80% and AIME24 reaching 46.7%, surpassing o1-preview - at a cost of \$42 versus thousands for baselines. Open-RS variants averaged 53.0%–56.3% on benchmarks, demonstrating RL’s viability for small LLMs. Releasing our code and datasets, we provide a framework for lightweight, reasoning-capable models, despite challenges like optimization stability, laying a foundation for future work.
@@ Appendix A Related Work
Reasoning in Large Language Models
A substantial body of research has investigated methods to enhance the reasoning capabilities and factual accuracy of large language models (LLMs). Early approaches predominantly relied on prompting techniques to elicit structured reasoning. For instance, scratchpad-style prompting encourages models to break down problems into intermediate steps (Nye et al. 2021), while verification mechanisms assess the correctness of generated outputs (Cobbe et al. 2021). Chain-of-thought (CoT) prompting has emerged as a particularly effective strategy, leveraging demonstrations of step-by-step reasoning to improve performance on complex tasks (Wei et al. 2022; Kojima et al. 2022; Reynolds and McDonell 2021). More recently, techniques such as intermediate self-reflection have been proposed to enable models to iteratively refine their reasoning processes (Shinn et al. 2023; Madaan et al. 2023).
In parallel, supervised fine-tuning (SFT) has been employed to embed reasoning capabilities directly into LLMs. Studies such as (Lewkowycz et al. 2022b) and (Rajani et al. 2019) demonstrate that fine-tuning on high-quality datasets can enhance problem-solving abilities. Notably, integrating CoT reasoning into SFT has shown significant promise; works like (Zelikman et al. 2022; Muennighoff et al. 2025; Ye et al. 2025) illustrate that fine-tuning with small, carefully curated datasets of CoT examples can yield substantial performance gains. However, these efforts have predominantly focused on large-scale LLMs, typically ranging from 7 billion to over 100 billion parameters. This reliance on massive models limits accessibility and practicality for resource-constrained settings, motivating the exploration of alternative approaches for smaller LLMs.
Reasoning with Reinforcement Learning
Reinforcement learning (RL) has emerged as a powerful paradigm for improving reasoning in LLMs, particularly for tackling complex, multi-step problems. Unlike SFT, which often optimizes for imitation of training data, RL enables models to learn from feedback, enhancing generalization to both in-distribution and out-of-distribution tasks (Chu et al. 2025; Yeo et al. 2025). Recent advancements underscore the efficacy of RL in this domain. For example, (OpenAI 2024b) and (DeepSeek-AI 2025) demonstrate that RL-based training can significantly boost reasoning performance, while (Team 2025a) explores scaling laws for RL-driven LLMs. These studies highlight RL’s ability to refine decision-making processes by optimizing for task-specific rewards, such as correctness or logical coherence.
@@ Appendix B Limitations & Discussion
Third, the multilingual nature of the base model, DeepSeek-R1-Distill-Qwen-1.5B, introduced unintended language drift after 150–200 steps, despite efforts to enforce English-only outputs via prompts in Experiment 3. This limitation reflects a trade-off in using a pre-trained, multilingual foundation, which, while efficient, complicates monolingual optimization. Finally, our evaluation focused exclusively on mathematical reasoning benchmarks, leaving the generalizability of our approach to other domains - such as scientific reasoning or coding - unexplored. These constraints highlight the need for cautious interpretation of our results within the specified scope.
Discussion
@@
Our findings reveal a nuanced trade-off between efficiency and reasoning depth in small LLMs. The rapid performance gains observed in the first 50–100 steps across all experiments (Insight 1) suggest that small, high-quality datasets can effectively bootstrap reasoning capabilities, aligning with prior work on data efficiency in RL (Chu et al. 2025). However, the subsequent degradation underscores a sensitivity to over-optimization under fixed length constraints, a challenge also noted in larger models like DeepSeek-R1 (DeepSeek-AI 2025). Experiment 2’s success with mixed difficulty levels (Insight 2) indicates that curriculum-like strategies could mitigate this. Meanwhile, the cosine reward’s stabilizing effect in Experiment 3 (Insight 3) suggests a promising direction for controlling reasoning verbosity, though it sacrifices peak accuracy compared to Experiment 2.
@@
Comparatively, our Open-RS variants achieved performance rivaling or exceeding state-of-the-art 1.5B models (e.g., DeepScaleR-1.5B-Preview) and even some 7B models, at a fraction of the cost and data volume. This efficiency challenges the prevailing reliance on massive datasets and computational resources in reasoning enhancement (OpenAI 2024b; Luo et al. 2025), offering a scalable alternative for resource-constrained environments. However, the persistent multilingual drift and length limitations point to inherent challenges in adapting multilingual base models and optimizing for complex tasks within tight bounds.
Future Directions
@@
These limitations suggest several research avenues. First, extending training duration or employing multi-stage length schedules could address truncation issues, allowing the model to handle harder problems without compromising stability. Second, incorporating a lightweight language reward or monolingual pre-filtering of the base model might mitigate language drift, enhancing output consistency. Third, expanding the benchmark suite to include non-mathematical domains would test the generalizability of our approach, aligning with broader AGI goals. Finally, exploring hybrid methods - such as combining GRPO with search algorithms like MCTS (Feng et al. 2024) - could further deepen reasoning capacity without significantly increasing resource demands.
@@
In conclusion, our work demonstrates that RL-based fine-tuning can unlock substantial reasoning potential in small LLMs, even under stringent constraints. By identifying key trade-offs and offering practical insights, we pave the way for developing efficient, reasoning-capable models that balance performance and accessibility - a critical step toward democratizing advanced AI technologies.
@@ Appendix C Datasets
Detail datasets are used in Section˜3.
  • AIME24444https://huggingface.co/datasets/AI-MO/aimo-validation-aime: 30 problems from the 2024 American Invitational Mathematics Examination, emphasizing advanced high-school-level reasoning.
@@
  • MATH-500 (Lightman et al. 2023b; Hendrycks et al. 2021): A subset of 500 problems from the MATH benchmark, sourced from various mathematics competitions and spanning algebra, calculus, and geometry.
  • Minerva (Lewkowycz et al. 2022a): 272 undergraduate-level problems across physics, biology, chemistry, economics, and other sciences, requiring quantitative reasoning (mathematics subset used).
  • OlympiadBench (He et al. 2024): 675 Olympiad-level problems in mathematics and physics, designed to challenge advanced reasoning abilities.
@@ Appendix D Baseline Models
The description of baseline models in Section˜3.
@@
    • Llama-3.1-70B-Instruct (AI 2024a): A 70B-parameter model optimized for instruction-following.
    • o1-preview (AI 2024b): A high-performing reasoning model from OpenAI.
@@
    • Qwen-2.5-Math-7B-Instruct (Yang et al. 2024): An RL-trained model with a reward model for mathematical reasoning.
    • rStar-Math-7B (Guan et al. 2025): Uses Monte Carlo Tree Search (MCTS) for deep reasoning and self-evolution.
    • Eurus-2-7B-PRIME (Cui et al. 2025): Employs the PRIME method with online RL and process rewards.
    • Qwen2.5-7B-SimpleRL (Zeng et al. 2025): Applies Proximal Policy Optimization (PPO) with rewards based on final answers.
@@
    • DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI 2025): The original untrained baseline.
    • DeepScaleR-1.5B-Preview (Luo et al. 2025): Fine-tuned with GRPO on 40,000 math problem-answer pairs across multiple RL stages.
    • Still-3-1.5B-Preview (Team 2025b): RL-trained with a focus on slow thinking (e.g., tree search) using 30,000 curated math examples.
@@ Appendix E Hyperparameter Setup
Table˜5 show parameters that used in training phase.